OpenStudy (anonymous):

Name the algebraic property demonstrated in the example below. (2 + x) + y = 2 + (x + y) Commutative Property of Addition Associative Property of Addition Distributive Property Transitive Property

OpenStudy (anonymous):

im thinking its C

OpenStudy (anonymous):

here, the GROUPING changed....

OpenStudy (anonymous):

so it's not C....

OpenStudy (anonymous):

is it d? since it channges?

OpenStudy (anonymous):

no.... transitive property is : if a=b and b=c, then a=c....

OpenStudy (anonymous):

uugh im getting stressed, so is it associative?

OpenStudy (anonymous):

HINT: grouping deals with how you associate.....

OpenStudy (anonymous):

here's a good way to remember these: commutative property deals with the ORDER of how things are operated on: a+b = b+a associative property deals with GROUPING things together and getting the same answer: (a+b)+c = a+(b+c)

OpenStudy (anonymous):

Solve the equation x − 3 = 8 11 5 −11 −5